Find the equation of the line through the given point that is perpendicular to the given line y=3x+5, Q(1,2)
Answer by jmg(22)

In the equation y = 3x + 5 , the number in front of x is the slope.
That means the slope is 3. The slope of a line perpendicular to that line would be the negative reciprocal which would be -1/3
Fill the slope along with the point Q into the point-slope form:
y - y sub 1 = m ( x - x sub 1) m stands for slope
y - 2 = (-1/3)(x - 1) distribute the -1/3
y - 2 = (-1/3)x + 1/3 bring the -2 across
y = (-1/3)x + 2 and 1/3